OpenStudy (kakay12):

Convert 550 cm into m. Please show your work.

OpenStudy (adonine):

5.5 m

OpenStudy (adonine):

thats the answer

OpenStudy (vijeya):

We know that 1m=100cm So what would 1 cm be equal to? It would be- 1 cm =1/100 m So here, 550 cm =550/100 m which will be equal to 5.5 m I hope that is helpful:)

OpenStudy (kakay12):

THANKS

OpenStudy (adonine):

medal